# A JWST Survey of the Supernova Remnant Cassiopeia A

### v1

@@ 8 Conclusions
We have uncovered a web-like network of unshocked ejecta filaments seen in [O iv] emission and resolved to 0.01 pc scales reflecting turbulent mixing processes and hydrodynamical instabilities that occurred shortly after core collapse of the progenitor star. We weakly detect emission possibly associated with [Fe viii] 5.446 m that had been predicted by Laming & Temim (2020), and estimate upwards of M of diffuse, unshocked Fe material. Many aspects of this distribution have significant relevance for core-collapse SN simulations, including a clear connection seen between the unshocked ejecta filaments and Chandra Fe–K emission. Correlations with NuSTAR are less clear but worth further investigation. This finding is also relevant for attempts to simulate light curves and spectra of unresolved extragalactic SNe, which must make assumptions regarding the fragmentation of the ejecta due to hydrodynamical instabilities. Ni-bubble expansion reduces the effective optical depth of non-radioactive material (Dessart et al., 2018; Dessart & Audit, 2019), and inappropriate treatment could mean that the current ejecta masses estimated for the majority of stripped-envelope SNe are systematically underestimated (Ergon & Fransson, 2022; Ergon et al., 2023).
@@
A large structure called the “Green Monster,” interior to the main shell and bright in MIRI images, is resolved for the first time. We conclude that this is most likely associated with a thick sheet (or sheets) of dust-dominated emission from shocked CSM seen in projection toward the remnant’s interior. The structure is pockmarked with small () round holes that have most likely been formed by high-velocity N- and He-rich ejecta knots that have pierced through the CSM and driven expanding tangential shocks. Support for this conclusion, including detailed analysis of MIRI/MRS and NIRSpec spectroscopy and correlated X-ray emission observed with Chandra, is provided in De Looze et al. (2024) and Vink et al. (2024). The discovery provides an exciting opportunity to probe mass loss of the progenitor system some years prior to explosion.
@@
Additional epochs of JWST observations would make it possible to improve characterization of the kinematic and elemental properties of the ejecta via proper motion measurements, while also permitting further study of nearby light echoes. Imaging mosaics with wider fields of view (reaching 5′ from the center of expansion versus the 3′ reach of our existing data) would increase opportunities for these measurements dramatically, and cover the NE/SW outflows to more accurately constrain their total mass, composition, and relationship to the original SN (Fesen & Milisavljevic, 2016). Additional spectroscopy, including widefield NIRSpec MSA observations across the entire SNR and targeted MIRI/MRS observations of regions interior to the main shell (especially ones that overlap with Ti emission), would be invaluable to improve estimates of chemical abundances and place the deepest limits possible for the total Fe yield.
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1947. Support for program #1947 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127.
D.M. thanks the Munich Institute for Astro-, Particle and BioPhysics for hosting the “Interacting Supernovae” program where interpretations used in this manuscript were developed. We are grateful for the Lorentz Center for hosting the workshop “Supernova Remnants in Complex Environments” where some ideas used in this manuscript were developed. Tracey DeLaney provided Spitzer IRS and VLA S-band data that were inspected as part of this paper. Brian Grefenstette provided the NuSTAR Ti map used in this paper. Una Hwang provided the Chandra 1 Msec image used in this paper. D.M. is grateful for conversations with Luc Dessart, who contributed ideas used in this paper. D.M. acknowledges invaluable support from Mark Linvill and Chris Orr who help manage computational resources used to complete this research.

### v2

@@ 8 Conclusions
We have uncovered a web-like network of unshocked ejecta filaments seen in [O IV] emission and resolved to 0.01 pc scales reflecting turbulent mixing processes and hydrodynamical instabilities that occurred shortly after core collapse of the progenitor star. We weakly detect emission possibly associated with [Fe VIII] 5.446 m that had been predicted by Laming & Temim (2020), and estimate upwards of M⊙ of diffuse, unshocked Fe material. Many aspects of this distribution are relevant for core-collapse SN simulations, including a clear connection seen between the unshocked ejecta filaments and Chandra Fe–K emission. Correlations with NuSTAR are less clear but worth further investigation. This finding is also relevant for attempts to simulate light curves and spectra of unresolved extragalactic SNe, which must make assumptions regarding the fragmentation of the ejecta due to hydrodynamical instabilities. Ni-bubble expansion reduces the effective optical depth of non-radioactive material (Dessart et al., 2018; Dessart & Audit, 2019), and inappropriate treatment could mean that the ejecta masses for the majority of stripped-envelope SNe are systematically underestimated (Ergon & Fransson, 2022; Ergon et al., 2023).
@@
A large structure called the “Green Monster,” interior to the main shell and bright in MIRI images, is resolved for the first time. We conclude that this is most likely associated with a thick sheet (or sheets) of dust-dominated emission from shocked CSM seen in projection toward the remnant’s interior. The structure is pockmarked with small () round holes that have most likely been formed by high-velocity N- and He-rich ejecta knots that have pierced through the CSM and driven expanding tangential shocks. Support for this conclusion, including detailed analysis of MIRI/MRS and NIRSpec spectroscopy and correlated X-ray emission observed with Chandra, is provided in De Looze et al. (in preparation) and Vink et al. (2024). The discovery provides an exciting opportunity to probe mass loss of the progenitor system some years prior to explosion.
@@
Additional epochs of JWST observations would make it possible to improve characterization of the kinematic and elemental properties of the ejecta via proper motion measurements, while also permitting further study of nearby light echoes. Imaging mosaics with wider fields of view (reaching 5′ from the center of expansion versus the 3′ reach of our existing data) would increase opportunities for these measurements dramatically, and cover the NE/SW outflows to more accurately constrain their total mass, composition, and relationship to the original SN (Fesen & Milisavljevic, 2016). Additional spectroscopy, including widefield NIRSpec mirco-shutter assembly observations across the entire SNR and targeted MIRI/MRS observations of regions interior to the main shell (especially ones that overlap with 44Ti emission), would be invaluable to improve estimates of chemical abundances and place the deepest limits possible for the total Fe yield.
We thank the referee for a very helpful reading of the Letter and suggestions that significantly improved its content and presentation. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #1947. Support for program #1947 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127.
This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311. We are grateful for the Lorentz Center for hosting the workshop “Supernova Remnants in Complex Environments” where some ideas used in this manuscript were developed. Tracey DeLaney provided Spitzer IRS and VLA S-band data that were inspected as part of this paper. Brian Grefenstette provided the NuSTAR 44Ti map used in this paper. Una Hwang provided the Chandra 1 Msec image used in this paper. D.M. is grateful for conversations with Luc Dessart, who contributed ideas used in this paper. D.M. acknowledges invaluable support from Mark Linvill and Chris Orr who help manage computational resources used to complete this research.
@@
The color composite images shown in the body of this paper are helpful for showing the relative differences between multiple filters of data, but they can also hide details that are present within the individual data sets, especially for our MIRI imaging. As indicated in Figure 2, some bandpasses are dominated by continuum emission and others contain both continuum and line emission. In Figures A1, A2, and A3 of this appendix, we show images of each filter of the NIRCam and MIRI imaging mosaics individually to allow these details to be assessed on a filter-by-filter basis.